OpenStudy (anonymous):

The theme of the cache is "Superheroes." Carefully graph the four lines with the following equations on the coordinate plane below to see the shape of the item. im going to post the equations in the replys

OpenStudy (anonymous):

this the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A. y = 2x + 13 B. 2x - y = 17 C. y =-1/2x + 13 D. x + 2y =-14

OpenStudy (anonymous):

can you put the cordenets for them

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(0, 13) (12, 7) (4, -9) (-8, -3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np